OpenStudy (anonymous):

The equation y=sqrtx, x=9, y=0, and a rectangle with its sides parallel to the axes and its left end at x=a. Find the dimensions of the rectangle having the maximum possible area.

OpenStudy (kirbykirby):

Are there any restrictions on \(a\) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (kirbykirby):

Try and relate the area of the rectangle with the \(\sqrt{x}=y\) function. You can see that:  So, the area of the rectangle is \(A=(9-a)(y)\) since y=sqrt(x) But, you know that at the vertical line x=a, you have the point \(y=\sqrt{a}\) So you get that \(A=(9-a)(\sqrt{a})\)